a car traveling 60 mph enters a highway and drives for 8 min,then a truck enters the highway going 68mph. How long will it take for the truck to pass the car
Answer by checkley77(12844)

D=RT
D=60T FOR THE CAR.
D=68(T-8/60)
SET THE 2 DISTANCES EQUAL & SOLVE FOR T
60T=68(T-4/30)
60T=68T-(68*4/30)
60T-68T=-272/30
-8T=-136/15
T=(-136/15)/-8
T=136/15*1/8
T=136/120
T=1.1333 HOURS OR 1 HOUR 8 MINUTES THE TRUCK WILL CATCH THE CAR.
PROOF:
60*1.1333=68(1.1333-.1333)
68=68